OpenStudy (anonymous):

Conner is 48 inches tall theres are 7 rides at the amusement park that conner can ride when he is 4 ft tall and 12 rides that he must be 5 ft. tall and 3 rides that be must be at least 3ft. tall to ride. What is the ratio of rides that Conner can ride now to the rides he cannot ride? A. 5:6 b. 7:15 c. 3:19 D. 10:21

OpenStudy (anonymous):

12 inches in a foot 48/12=4ft he can ride 10 rides he can ride and 12 rides he cannot this makes the ratio \[\frac{5}{6}\]

OpenStudy (anonymous):

This question is horrible. Why in the world would they put 3 feet last =[

OpenStudy (anonymous):

total rides = 7+12+3=22....5ft = 12 so 10:12 in can vs can't which comes to 5:6

OpenStudy (anonymous):

Can you please Aoutkas give the medal to paulk thanks to boths